OpenStudy (watchmath):

All Calculus students dream :) Find all pair of non constant functions f, g such that (fg)'=f'g' and (fg)''=f''g''

OpenStudy (anonymous):

The most obvious solution (but not the one you're looking for) \[\text{Let } f = {1\over g}\]

OpenStudy (watchmath):

well I don't think that works!

OpenStudy (anonymous):

Oh. true.

OpenStudy (anonymous):

hey watchmath, i got disconnected earlier

OpenStudy (anonymous):

the program just froze in the middle of convo

OpenStudy (anonymous):

ok lets see, use product rule

OpenStudy (anonymous):

(fg)'= f'g' and (fg)''=f''g'' but (uv)' = u'v + uv' , , so ...

OpenStudy (anonymous):

one sec , dont give answer

OpenStudy (anonymous):

i suspect theres an e^x in there

OpenStudy (anonymous):

(fg)' = f' g + f g' = f' g'

OpenStudy (anonymous):

(fg) ' = f' (g / g' ) + f/f'

OpenStudy (anonymous):

since i divided both sides by g ;

OpenStudy (anonymous):

errr, (fg)' = f' g + f g' = f' g' f ' = f ' ( g/g') + f

OpenStudy (anonymous):

so f = f' ( 1 - g / g ' )

OpenStudy (anonymous):

nevermind, lets use the second assumption

OpenStudy (anonymous):

no hints, one sec

OpenStudy (anonymous):

(fg)'' = ( f ' g' ) ' = f '' g ' + f ' g''

OpenStudy (anonymous):

so we have two equations fg ' + f ' g - f ' g' = 0 f '' g' + f' g'' - f'' g'' = 0

OpenStudy (anonymous):

integral f '' *g'' = (fg)' , integral (fg)' = fg

OpenStudy (anonymous):

lets integrate by parts

OpenStudy (anonymous):

ok i need a hint, im stumped

OpenStudy (watchmath):

consider fg ' + f ' g - f ' g' = 0 as a differential equation in g :)

OpenStudy (anonymous):

WAIT, FROM THE FIRST equation*

OpenStudy (anonymous):

g/g' + f/f' = 1, g'/g'' + f' / f'' = 1

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

what kind of diferential equation is it

OpenStudy (watchmath):

It is linear

OpenStudy (anonymous):

ok so ( f - f') g ' + f' g = 0

OpenStudy (anonymous):

i dont know how to solve that, only if they are constant, but lets assume they are

OpenStudy (anonymous):

then (f-f')^2 + f' = 0

OpenStudy (watchmath):

Well that equation can be made into something of the type: y'+p(t)y=0

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i dont think there will be any nonconstant functions satisfying the given criteria....but still searching for a proof. By the way the the problem is good!

OpenStudy (watchmath):

There are! If you just want to find some you can guess one! What kind of functions are the obvious guess?

OpenStudy (anonymous):

ok i have f' ( g/g' -1 ) + f - 0

OpenStudy (anonymous):

f' ( g/g' -1 ) + f = 0

OpenStudy (anonymous):

see g=(some constant)*g' or same for f...doesnt work...there could be more complex functions than e^kx...what are ur guesses watchmath?

OpenStudy (watchmath):

Here is one of the example \(f=e^{3x/2},g=e^{3x}\)

OpenStudy (watchmath):

Now I believe you can easily characterize which eponential functions \(f=Ae^{Bx},g=Ce^{Dx}\) satisfy the requirement.

OpenStudy (anonymous):

bye